### v1

# Semantic Tree Inference on Text Corpa using a Nested Density Approach together with Large Language Model Embeddings

Abstract
Semantic text classification has undergone significant advances in recent years due to the rise of large language models (LLMs) and their high dimensional embeddings. While LLM-embeddings are frequently used to store and retrieve text by semantic similarity in vector databases, the global structure semantic relationships in text corpora often remains opaque.
Herein we propose a nested density clustering approach, to infer hierarchical trees of semantically related texts. The method starts by identifying texts of strong semantic similarity as it searches for dense clusters in LLM embedding space. As the density criterion is gradually relaxed, these dense clusters merge into more diffuse clusters, until the whole dataset is represented by a single cluster - the root of the tree. By embedding dense clusters into increasingly diffuse ones, we construct a tree structure that captures hierarchical semantic relationships among texts.
We outline how this approach can be used to classify textual data for abstracts of scientific abstracts as a case study. This enables the data-driven discovery research areas and their subfields without predefined categories. To evaluate the general applicability of the method, we further apply it to established benchmark datasets such as the 20 Newsgroups and IMDB 50k Movie Reviews, demonstrating its robustness across domains.
Finally we discuss possible applications on scientometrics, topic evolution, highlighting how nested density trees can reveal semantic structure and evolution in textual datasets.
1 Background
Text classifications date back at least to bag of words models [bagofwords]. In these early approaches, each text is represented as a vector of word frequencies, disregarding the order of words. However, word order is a key semantic feature as changing it can completely alter the meaning of a phrase.
Subsequent refinements introduced term frequency and inverse document frequency (tf-idf) normalization [tfidf], which weights words by their relative importance within a corpus. Later, Latent Semantic Analysis (LSA) [lsa-first, lsa, lsa-intro] applied singlar value decomposition to capture latent relationships among words and documents, enabling more robust document clustering and topic modeling. Easier to scale and based on probabilistic properties instead of linear algebra Latent Dirichlet Allocation (LDA) [lda] and Hirachical Dirichlet Allocation (HDA) [hda] have overtaken the field of topic modelling. Neural network based approaches such as word2vec [word2vec] have arrived in the middle of the last decade. These models could learn the relationships between words in a semantic proximity sense, but still not in a grammatical sequence. The rise of the transformer model [transformer] and the popularization of LLMs allowed for the use of LLM embeddings for document classification, and document retrieval. The transformer model is superieur as it efficiently encodes and uses the position of the words in the text as it generates embeddings. LLM embeddings have a such a vast field of applications and [llm-embed-chatbot, llm-embed-clust] commonly found in retrieval augmented generation (RAG) pipelines [rag, llm-embed-rag]. All of the aforementioned methods project text into a mathematical embedding space, where pairwise distances and clustering can be applied to reveal latent semantic structures. In parallel, hierarchical density clustering techniques have proven highly effective in fields such as biological species classification, phylogenetics [mnhn-tree-tools], and molecular dynamics, where they are used to analyze protein conformations and structural evolution [md-tree-tools].
Here, we extend these hierarchical density clustering methods to textual embeddings, constructing tree-based representations of semantically related texts. This approach goes beyond flat clustering by revealing how topics, subjects, or emotions (depending on the embedding model) relate to one another within a coherent hierarchical framework.
2 Introduction
Topic modeling and semantic clustering have long aimed to uncover the latent structure of large text corpora. With the rise of large language models (LLMs) and their embeddings, it has become possible to represent textual semantics with unprecedented precision. However, most embedding-based clustering approaches produce flat partitions of the semantic space, providing limited insight into how topics or concepts are hierarchically related or evolve across scales.
In this work, we introduce a hierarchical density-based clustering approach that infers tree structures from LLM embedding spaces. The method builds upon previous applications of hierarchical density clustering in bioinformatics and molecular dynamics, where it has been successfully used to construct phylogenetic or conformational trees from complex, high-dimensional data [mnhn-tree-tools, md-tree-tools]. By adapting this framework to text embeddings, we enable the reconstruction of semantic hierarchies. We show how dense clusters of highly similar documents merge into broader, more diffuse clusters as the similarity threshold is relaxed.
To construct such hierarchical structures, we employ a nested density-based clustering procedure inspired by the DBSCAN algorithm [dbscan]. The core idea is to identify compact, high-density regions of semantically similar embeddings and then progressively relax the density criterion to reveal broader, more diffuse groupings. This iterative relaxation enables clusters discovered at a high-density level to merge into larger clusters at lower density levels, naturally producing a hierarchical tree. Formally, clusters of semantically similar texts that that are densly connected in their respective embedding space are determined by the condition:
$\rho_{L}>\frac{\mathrm{minpts}}{V(\epsilon_{L})},$ (1)
where $\rho_{L}$ represents the denstiy at a given iteration $L$. The right hand side is given by starting parameters to the well known Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [dbscan], with $\epsilon_{L}$ defining a radius at iteration $L$ and $\mathrm{minpts}$ a minimum number of points within this radius, effectively defining the density of a continious dense ensemble to be found by the algorithm. To explore the semantic structure at multiple scales, the radius $\epsilon_{L}$ is incrementally expanded by a small step $\Delta\epsilon$ in each iteration. This results in a sequence of density thresholds: As $\epsilon$ increases, the volume in equation (1) expands leading to lower densities and the identification of more diffuse clusters in each iteration. These new clusters encapsulate denser ones from previous iterations, allowing us to embed fine-grained semantic clusters into coarser ones, thereby forming a hierarchical tree of semantic relationships. For a subsequent layer (iteration) of the tree $L+1$ it follows, equation (2):
$\rho_{L+1}>\frac{\mathrm{minpts}}{V(\epsilon_{L}+\Delta\epsilon)},$ (2)
or alternatively $\epsilon_{L}=\epsilon_{L-1}+\Delta\epsilon$. In our previous work [mnhn-tree-tools], we demonstrated that using this algorithm allows us to construct a hierarchical tree structure by embedding dense clusters into progressively more diffuse clusters until all elements of the conformational space merge into a single cluster - the root of the tree. Figure 1 outlines the tree building procedure conceptually. On the left, a text embedding space is shown, where each colored sphere represents a piece of text. On the right the tree built by the algorithm, from the text embedding space on the left, is shown. One can observe how the grape, and orange colored texts fusion as the density criteria is relaxed into the blue ensemble, encompassing both clusters. At an intermediate step the green cluster stays separate, while the violet texts are not encompassed and seen as outliers. Finally, as the density criterion is further relaxed all texts are only part of one cluster, represented in gray on the right tree-outline.
Figure 1: Tree building from a dataset represented by its embedding space. On the left, an embedding space with clusters of different densities is shown. One can see that the orange, and grape colored clusters embed into a larger blue ensemble. A conceptual tree built from such a two dimensional embedding space is shown on the right.
Our algorithm allows an organization for instance to show how documents within the organization are related. It initially finds separated clusters of highly semantically similar documents (the outer leafs of the tree), that merge, layer by layer together to more diffuse clusters, until all documents form a single cluster (the root of the tree).
We demonstrate our approach on several corpora, including collections of scientific abstracts and benchmark datasets such as 20 Newsgroups and IMDB 50K reviews. In these experiments, we show how the inferred semantic trees capture both fine-grained clusters of related documents and higher-level thematic organization. Applied to the publication records of institutions such as TU Wien and the American University of Beirut, our method reveals naturally emerging research areas and structural relationships between topics.
Once a semantic tree is constructed, an important challenge lies in interpreting and annotating its clusters. We address this by leveraging LLMs themselves: by prompting a model with the text content of each cluster, we can automatically generate concise, human-readable summaries or titles for each node in the hierarchy. This creates an interpretable and fully automated pipeline for semantic structure discovery in large text corpora.
3 Materials and Methods
3.1 Datasets
The primary goal of this study was to develop a pipeline for the automated classification of research fields based on textual similarity. While motivated by this scientific application, the proposed approach is general and can be applied to a broad range of text corpora. To ensure linguistic consistency and to avoid potential errors caused by multilingual embeddings and the tree building process, which would in such a case probably attributed different branches to different languages, we restricted our analysis to texts written in English.
We selected the following datasets for our study:
  • AUB: English abstracts of publications authored by the American University of Beirut (AUB) from 2018 to October 2025. The data were provided by the AUB University Libraries. To ensure comparability, the roughly the same temporal window and language filters, only english abstracts, as for the TU Wien corpus were applied.
  • TU Wien: English abstracts of publications affiliated with TU Wien between 2018 and April 2025. The data were provided by the TU Wien Service Unit for Research Information Systems as aggregated extracts from the institutional repository (reposiTUm), Scopus, and Dimensions databases. Duplicates were removed, and only records containing English abstracts were retained.
  • 20 Newsgroups: The 20 Newsgroups dataset is a widely used benchmark corpus for text classification and topic modeling. It consists of approximately 20,000 newsgroup posts collected from 20 distinct discussion groups on topics such as science, politics, sports, religion, and technology [20newsgroups].
  • IMDB 50K Reviews: A corpus of 50,000 movie reviews from the Internet Movie Database (IMDB), each labeled with a binary sentiment (positive or negative) [50kmovies].
  • AG News: A large-scale news dataset consisting of short article titles and descriptions categorized into four topics: World, Sports, Business, and Science/Technology. [agnews].
  • Theses-fr-en: English abstracts of theses defended in France: The French government publishes by means of its open data initiative a dataset containing all theses defended in France since 1985. We filtered this dataset for english abstracts. This data was obtained from the data.gouv.fr platform.
Statistics about the different datasets are shown in table 1.
n(Texts) mean stddev
TU Wien 27506 168.85 73.79
AUB 15982 237.73 98.24
20newsgroups 11314 287.47 541.46
50k-movies 50000 231.16 171.34
ag_news 127000 37.84 10.09
theses-fr-en 180939 261.89 112.70
Table 1: Statistics about the text corpa used. n(Texts) is the number of unique texts in the dataset, mean and stddev represents the mean and standard deviation of the number of words contained in each text of the dataset.
3.2 Embeddings
Embeddings for the datasets outlined in section 3.1 were generated using either the Qwen3-Embedding-8B [qwen3embedding] or SFR-Embedding-Mistral [SFRembedding] model. These models were chosen selected based on their high performance in the MTEB-Leaderboard found on Huggingface, and the their practical deployability on the computing infrastructure accessible to us.
4096 dimensonal embedding vectors obtained for each text from a dataset were stored in an in house developed vector database.
Each text sample was transformed into a 4096-dimensional embedding vector. The resulting embeddings were stored in an in-house developed vector database optimized for high-throughput semantic similarity retrieval.
3.3 Tree Building
3.3.1 Pairwise Distance and Dimensionality Reduction
The pairwise distance between embeddings serves as the foundation for the hierarchical tree construction process. We employed the cosine distance on the full-length embedding vectors as well as the $L_{2}$-norm on reduced-dimensional representations obtained via Principal Component Analysis (PCA).
Dimensionality reduction proved crucial, as PCA functions as a feature selection mechanism that prioritizes components with maximum variance. Empirically, we observed that reducing the number of components leads to more clearly separated and interpretable cluster structures. While the original embedding vectors contain 4096 dimensions, spectral analysis indicated significant redundancy among components. The choice of the number of principal components thus directly affects the resolution and spread of the resulting semantic tree.
3.3.2 Adaptive Density Clustering with DBSCAN
For each dataset, we constructed trees using an adaptive density-based clustering procedure built upon successive runs of the DBSCAN algorithm. Each DBSCAN iteration identifies clusters at a specific density level, controlled by a triplet of parameters: the initial neighborhood radius $\epsilon_{0}$, the incremental radius step $\Delta\epsilon$, and the minimum number of points within the radius ($\mathrm{minpts}$).
In each successive run, the radius is increased by $\Delta\epsilon$ while keeping $\mathrm{minpts}$ constant. This progressively lowers the effective density threshold and allows previously dense clusters to merge into more diffuse ones. Whenever the number of clusters found in a given iteration is smaller than in the preceding step, the newly identified clusters are stored, and the clusters from the prior iteration are embedded within them. This iterative embedding process builds the hierarchical tree structure described in Figure 1, where dense clusters at outer layers merge stepwise into broader, semantically diffuse clusters toward the root.
3.4 Tree Annotation
The resulting trees consist of multiple hierarchical layers. Each layer represents a set of clusters that form the nodes at that level, while edges connect clusters of higher density (outer layers) to broader, lower-density clusters (inner layers). This structure allows fine-grained semantic groups to merge progressively into coarser domains.
To realize our goal of a data-driven research field classification pipeline, we sought to automatically annotate each node with a representative topic label. The underlying assumption is that smaller, high-density clusters correspond to specialized research areas, while broader clusters capture more general scientific domains.
To generate these annotations, we employed a Large Language Model (LLM) capable of handling large textual contexts. Specifically, we used the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4] in 6-bit quantized form, which provides a nominal context window of up to 11 million tokens. Due to technical constraints, our inference pipeline successfully operated with a maximum context size of approximately 600,000 tokens.
Both the TU Wien and AUB corpora were fully annotated, encompassing all English-language scientific abstracts published since 2018. For each node, the LLM received the text of all abstracts within that cluster and was prompted to infer a unifying research field label.
[…]
In the case were the cluster had to be sampled because it contained more than 1500 abstracts all resulting labels were subsequently fed into a second LLM pass. Again the gpt-oss-20b model [gpt-oss] in 8-bit quantized form is used in such a case where the model inferred the overarching research field encompassing all $s$ subfields. An overview about the different paths that annotation might undergo is outlined in Figure 2.
Figure 2: Shematic represantation of the tree node / cluster annotation process
Tree annotation is computationally intensive. The AUB tree required annotating 33,862 text files, while the TU Wien tree required 62,012, including oversampling. Inference was performed using the llama.cpp [llama-cpp] engine on the MUSICA supercomputer (Austria). Up to 80 NVIDIA H100 GPUs were employed concurrently during annotation. Each compute node with four H100 GPUs (80 GB VRAM each) handled batches of approximately 600,000 tokens per inference instance. Multi-node inference, which could allow for a higher context token limit, was not attempted as MUSICA is currently in its test phase; instead, parallelism was achieved by distributing text files across multiple nodes for simultaneous inference.
3.5 Label Transcription
The theses-fr-en dataset is annotated and contains a research field (discipline, in French) for each abstract. However, these fields can in many cases be freely chosen by the thesis author, resulting in 13,599 unique discipline labels. In a first test we cross annotated these unique discipline labels into a binary classification of either humanities or natural sciences. We then proceeded to obtain a more coherent and standardized annotation, we used the gpt-oss-20b model to classify all 13,599 labels according to the OECD Fields of Science (FOS) 2002 classification [oecd-fos]. The resulting transcribed labels were then used to color the trees inferred from this dataset, allowing us to assess whether specific scientific fields tend to separate into distinct and meaningful branches.
3.6 Tree Coloration and Visualization
To facilitate interpretation and qualitative validation of the inferred trees, we applied coloration schemes that project existing labels or metadata onto the tree structure. This process enables us to visualize how known categories or attributes distribute across different branches of the tree.
For example, in the case of the 20 Newsgroups dataset, tree coloration reveals how posts from individual newsgroups are positioned within the hierarchical semantic structure. Similarly, for the TU Wien dataset, coloration based on existing research classifications allowed us to assess how well our hierarchical density clustering recovers semantically coherent domains.
Tree coloration was performed using utilities provided by MNHN-Tree-Tools [mnhn-tree-tools]. In particular, we employed the tree_map_for_split_set tool to map predefined clusters or labeled subsets onto the corresponding nodes of the hierarchical tree. The resulting color maps were generated for visualization using the newick-utilities package [newick-utilities], a software suite for the manipulation and visualization of phylogenetic trees.
This approach allowed us to produce colored tree visualizations for each dataset. In the 20 Newsgroups case, individual trees were colored according to each newsgroup category, highlighting topic segregation and overlap. For the IMDB 50K reviews dataset, we colored the tree by sentiment polarity, enabling a visual inspection of how positive and negative reviews cluster across different density layers.
For the annotated datasets from TU Wien and AUB, we required a high-resolution visualization that allows all node-level labels to be displayed and inspected. The resulting tree images are extremely large, making standard static rendering impractical. To enable detailed exploration, we developed a tile-based viewer application with two coordinated windows. One window displays the full, non-annotated tree, while the second window shows a zoomed view of a selected region, including the labels for the nodes currently in focus. Navigation is performed either by clicking on the full-tree view, which updates the zoomed window to the selected location, or by dragging within the zoomed window, which moves a red cursor on the full-tree view to indicate the corresponding position.
4 Results
4.1 Efficiency of PCA and the $L_{2}$-norm Compared to Cosine Distance
Cosine distance is widely used for comparing high-dimensional embedding vectors and is the default similarity measure in many retrieval-based applications, including RAG pipelines. For this reason, we initially expected cosine distance to be the natural choice for the DBSCAN iterations underlying our tree construction algorithm. However, trees constructed directly from full 4096-dimensional cosine distances consistently exhibited limited structural resolution: branches were poorly separated, fine-grained clusters failed to emerge, and the overall tree topology did not yield significant insights.
In contrast, applying Principal Component Analysis (PCA) prior to clustering dramatically improved the quality and expressiveness of the resulting trees. Experiments with 10, 5, and 2 principal components showed that lower-dimensional PCA representations consistently produced more “spread out” trees with clearer branching and more detectable fine-scale semantic substructure. We suggest that this effect is due to the inherent ability of PCA to yield principal components that maximize variance and that this feature selection process has a positive effect on our clustering algorithm. A similar effect was noted previously in the algorithmic supplement of [mnhn-tree-tools].
A comparison between trees obtained with cosine distance and with the $L_{2}$-norm applied to PCA-reduced embeddings is shown in Figure 3. From left to right, the trees become more hierarchically detailed and reveal progressively finer semantic subclusters. For these illustrative examples, the parameters were chosen using a rudimentary bisection search to produce visually optimized trees. The parameter settings were as follows: for cosine distance, initial $\epsilon=0.275$, $\Delta\epsilon=0.001$, and $\mathrm{minpts}=5$; for the $L_{2}$-norm on 10 principal components, initial $\epsilon=0.091$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$; and for the $L_{2}$-norm on 2 principal components, initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$. The trees shown in 3 were built from embeddings obtained form the Qwen3-Embedding-8B model.
Figure 3: A comparison of trees built from the 20 Newsgroups dataset under different distance measures
4.2 Detection of Topics and Semantics
Depending on the dataset, we applied different evaluation strategies to assess whether, and to what extent, the method proposed in this work produces meaningful and reliable results, given the computational constraints of our setup. The following subsections outline these evaluations and illustrate the behavior of our hierarchical density-based tree construction across diverse text corpora.
4.2.1 20 Newsgroups
To evaluate whether our tree construction method recovers meaningful semantic structure, we applied it to the 20 Newsgroups dataset. The resulting trees were colored according to posting labels, as shown in Figure 4. For each newsgroup we generated a separate color map: posts from the target newsgroup are shown in blue, and posts from all other newsgroups in orange. Both color intensities follow a logarithmic scale proportional to the number of posts in each cluster, with white indicating the absence of posts of a given type.
The inferred tree, as highlighted in figure 4, recovers many interesting relationships between the newsgroups:
  • Sports-related groups. The groups rec.sport.baseball and rec.sport.hockey form a shared subtree located close to the root of the full tree. This indicates strong internal similarity between the two sports discussions, combined with substantial separation from most other topics.
  • Automotive-related groups. The groups rec.autos and rec.motorcycles cluster tightly in a small dedicated sub-branch. A fraction of their posts appears in regions of the tree that correspond to scientific or technical discussions, likely reflecting threads involving engineering aspects of vehicles.
  • Religion and politics. The groups soc.religion.christian, alt.atheism, and talk.religion.misc form a joint subtree that also contains the political groups talk.politics.guns, talk.politics.mideast, and talk.politics.misc. This intersection reflects the topical proximity and overlapping discussions between religion and politics within the dataset.
  • Computer-related groups. The comp.* groups, including comp.sys.mac.hardware, comp.windows.x, comp.graphics, and comp.os.ms-windows, form a coherent subtree corresponding to computer hardware and software discussions. Some overlap with sci.electronics is visible, consistent with semantically similar content.
  • Marketplace group. The misc.forsale posts distribute across multiple regions of the tree, often co-occurring with the automotive and computer subtrees, suggesting that sale-related messages tend to mirror the underlying topical focus of corresponding technical groups.
  • The cryptography group. The sci.crypt group appears in a distinct set of branches close to the root, and is clearly separated from both the scientific and computer-related subtrees. This separation may stem from the unique vocabulary and specialized discourse typical of cryptography discussions.
Overall the structure seems to intuitively make sense, outlining intervoven fields of discussion. Importantly, our toolchain does not merely cluster newsgroups into isolated groups; it also conveys the relative proximity and semantic distance between them. The resulting tree therefore serves as a compact visual instrument that captures both fine-grained cluster structure and the broader topology of topic similarity within the dataset, all within a single interpretable representation.
Figure 4: Representations of the tree obtained from the 20 Newsgroups dataset. Trees were built from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] embeddings. Parameters used to build tree were: initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
4.2.2 IMDB 50K Reviews
The IMDB 50K Reviews dataset was used to evaluate whether our tree construction method, together with the underlying embedding models, is able to separate reviews by emotional polarity. We therefore built a tree from this dataset following the same procedure as for the other corpora. The resulting tree was then colored analogously to the 20 Newsgroups experiments: clusters containing predominantly positive reviews were assigned a blue shading, while clusters containing predominantly negative reviews were shaded in orange. As before, color intensity follows a logarithmic scale, with white indicating the absence of reviews of a given label and deeper blue or orange indicating higher concentration. Overlaps appear in darker, greenish to black tones.
The resulting tree is shown in Figure 5. Compared to the other datasets, the IMDB tree appears noticeably less structured, suggesting that the underlying corpus is more homogeneous and lacks strong semantic subdivisions beyond sentiment. Nonetheless, we observe a few small subtrees near the periphery of the tree in which positive reviews form a clear majority. Additional localized regions with a majority of positive or negative reviews can also be identified. We speculate that these structures arise from repetitive linguistic patterns that occur more frequently in either positive or negative reviews. However, they do not form large coherent branches, and overall the dataset does not separate cleanly by sentiment.
This observation suggests that the embedding models used: Qwen3-Embedding-8B [qwen3embedding] and SFR-Embedding-Mistral [SFRembedding] encode emotional polarity only to a limited extent. The results highlight the importance of choosing embedding models appropriate for the expected semantic structure when constructing trees.
We also attempted to annotate the IMDB tree with the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4]. The goal was to obtain genre-level descriptions of tree regions, under the assumption that certain movie genres might correlate with predominantly positive or negative reviews. This attempt was unsuccessful: the model did not produce reliable or meaningful genre assignments, and more fundamentally, such information appears to be largely absent from the review text itself. As such, deriving movie genres from this dataset seems infeasible with current methods.
Figure 5: The IMDB 50K Reviews dataset colored by sentiments: The tree was built using the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.5$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$.
4.2.3 AG News
We further built trees from the AG News dataset and colored them according to the four available labels: World, Sports, Business, and Science/Technology. For each label, a separate color map was generated. Items belonging to the target label are highlighted in blue, while all other items are shown in orange. As in previous visualizations, we use a logarithmic color gradient to represent the number of articles in each cluster, ranging from white (no articles of that type) to fully saturated colors (all articles in the cluster belong to the target label). The resulting trees are shown in Figure 6.
Figure 6: The tree obtained from the AG News dataset in different colorations. The tree shown was inferred from the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings. Parameters yield to the tree building algorithm were: initial $\epsilon=0.225$, $\Delta\epsilon=0.0005$, and $\mathrm{minpts}=5$
We observe that the world and sports categories form clearly separable structures: both branch off early from the root of the tree, resulting in two well-defined subtrees with only minimal contamination from other labels. Around the root, however, a small number of scattered articles appear, suggesting that certain headlines are not easily classifiable by the embedding models. These headlines seem to be very distant from the other news items, only joining the root at a very diffuse density.
The remaining two categories, business and science/technology, show substantial overlap. They jointly occupy the third major branch of the tree, and only small terminal sub-branches demonstrate consistent label-specific structure. We attribute this to the general type of embedding models that we are using here and that business headlines frequently mention science and technology and vice versa in this dataset.
4.2.4 The French Theses Dataset: Theses-fr-en
To further evaluate whether our method can identify and organize broad and fine-grained disciplinary structure, we applied our tree-building pipeline to the Theses-fr-en dataset, which consists of English-language abstracts of French doctoral theses. The dataset contains author-supplied disciplinary annotations, but these labels are highly heterogeneous and unstandardized, yielding 13,599 unique descriptors. Therefore, as outlined in Section 3.5, we first performed a cross-annotation step to obtain more tangible labels.
As an initial experiment, we mapped the original annotations to a binary classification: humanities versus natural sciences. The resulting tree clearly separated humanities theses from the remainder of the corpus, forming a large, coherent branch as shown in Figure 7A. Encouraged by this result, we proceeded to apply a more detailed transcription based on the OECD Fields of Science and Technology (FOS) 2002 classification scheme [oecd-fos].
In the tree three major branches are easily identifyable. The branch identified as “humanities” in the binary test contains predominantly Social Sciences (5) and Humanities (6) according to the FOS taxonomy. Social sciences, however, extend beyond this primary branch, with small isolated scattered clusters appearing throughout the tree.
The remaining two major branches contain the Natural Sciences (1), but with different structural emphases. Strikingly, Engineering and Technology (2) appears almost exclusively within the very large branch that occupies nearly the entire upper half of the tree outlined in Figure 7. In contrast, Medical and Health Sciences (3) predominantly occupy the smaller of the two natural-science branches, located on the left side of the tree. Examination of FOS subclasses shows that this medical branch also contains theses classified as Biological Sciences (1.6), reflecting their conceptual proximity.
The pattern of overlaps among FOS categories, particularly the fact that Natural Sciences (1) span both the engineering-dominant and the medical-dominant branches, highlights structural ambiguities inherent in the current FOS hierarchy. Although (1), (2), and (3) are defined as top-level categories, the tree suggests that Engineering and Technology (2) and Medical and Health Sciences (3) are more closely related to different subregions of the Natural Sciences (1) than to each other.
Moreover, the tree indicates that (2) and (3) merge into a common branch earlier than Social Sciences (5) and Humanities (6). After the social-science and humanities subtree, numerous small and weakly structured clusters enter the main trunk of the tree (lower right quadrant), indicating a sizeable number of theses that form islands in the embedding space and are not well represented by broad disciplinary groupings.
All three major branches further subdivide into smaller subtrees toward their terminal regions. However, we did not find consistent evidence that these subtrees correspond cleanly to the FOS subclasses. This lack of correspondence suggests that the tree structure could support a data-driven reorganizing of disciplinary taxonomies, potentially capturing conceptual relationships that are absent, or obscured, in existing classification systems. This observation directly motivates the subsequent section of the paper.
Figure 7: Tree generated from the French Theses theses-fr-en dataset using the first two principal components obtained from the embeddings yield by the SFR-Embedding-Mistral [SFRembedding] model:
A: Coloration according to binary classification Humanities and Natural Sciences.
B: Coloration according to the main groups of the OECD FOS classification.
C: Highlight of the subgroups of the Natural Sciences group according to OECD FOS.
Parameters for tree building were: initial $\epsilon=0.23$, $\Delta\epsilon=0.001$, $\mathrm{minpts}=5$.
4.2.5 Annotated Institutional Trees: TU Wien and AUB
Our analysis of the two institutional datasets (TU Wien and AUB) differs from the benchmark datasets presented earlier. In the benchmark cases, we projected known labels onto the inferred trees to assess how well embeddings recover an existing classification. In contrast, the institutional datasets are used to examine research structures ab initio: rather than testing whether a predefined taxonomy is recovered, we investigate how the tree itself organizes the institutional research landscape. Although the embedding models were taken trained on external corpora, and not further fine tuned, the resulting tree structure is independent of any curated ontology. This allows us to explore institution-specific disciplinary profiles, structural proximities between fields, and potential emerging clusters or cross-disciplinary connections. The results demonstrate that our method can capture meaningful organizational patterns without relying on predefined categories.
  • At the American University of Beirut, medicine and health-related disciplines dominate the research corpus, reflecting the strong presence of the university hospital. This dominance is visible in the tree structure, where the entire upper half of the circular tree forms the largest and most cohesive subtree. Within this region, several well-defined subbranches correspond to specific healthcare domains, including health and nutrition, health policy and systems research, and public health and epidemiology.
    A distinct humanities subtree is also present, characterized by numerous subbranches labeled as Middle Eastern Studies and Conflict Resolution. In addition, an engineering subtree that covers engineering, physics, and mathematics, is clearly visible. Notably, this engineering subtree merges with the humanities subtree at a higher density than with the medical subtree, suggesting a closer structural relationship between engineering and humanities research topics within this corpus than might be expected, which might be characteristic of the university’s liberal arts tradition.
    To the left of these major subtrees, smaller clusters related to both engineering and humanities fields branch into the tree and eventually merge with the broader medical subtree. Beyond this point, several near-linear terminal branches from a wide range of disciplines connect to the tree as outliers before joining the main structure.
    The full tree is shown in Figure 8 and can be explored in detail using our interactive tree viewer at https://genealogy.sematlas.com/. As in the other figures, color intensity is logarithmically scaled from white (no articles in the branch) to black (all articles in the dataset).
  • Technische Universität Wien (TU Wien)
    The tree structure for TU Wien differs markedly from that of AUB. As a technical university, TU Wien has a more focused research portfolio, which results in a tree that appears more homogeneous overall. Nevertheless, the tree reveals several large and meaningful subtrees that highlight the institution’s disciplinary strengths.
    At the upper right of the tree, we observe a subtree corresponding to quantum mechanics, quantum optics, and quantum-scale materials. This region merges smoothly into a broad materials science subtree, which occupies much of the upper left quadrant of the circular tree.
    Also in the upper left quadrant is a large subtree corresponding to information technology and computer science, which subdivides into numerous smaller branches. Further branches that merge into the main tree at later stages correspond to nanotechnology and semiconductor engineering. Additional clusters represent infrastructure planning and environmental engineering. One of the earliest large branchings near the root corresponds to mathematics and theoretical physics, forming a distinct foundational science substructure.
    The TU Wien tree is shown in Figure 9, and an interactive exploration is available through our tree viewer at https://genealogy.sematlas.com/. As with the AUB tree, color intensity follows a logarithmic scale from white (no articles) to black (all articles in the dataset).
Figure 8: Tree generated from the AUB dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0045$, $\Delta\epsilon=0.00001$ and $\mathrm{minpts}=5$.
Figure 9: Tree generated from the TU Wien dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings. The tree was built with the following parameters: initial $\epsilon=0.0028$, $\Delta\epsilon=0.000001$ and $\mathrm{minpts}=5$.
We note, as can also be verified using our custom tree viewer, that the automatic annotation of clusters is not without limitations. In many cases, the LLM assigns overly general labels—such as engineering—to large or heterogeneous clusters. Moreover, clusters that are structurally identical across different annotation runs sometimes receive different labels, suggesting that the model struggles with processing extremely large input contexts and exhibits instability in its naming decisions. Despite these limitations, the annotation process provided substantial insights into both institutions, their publication profiles, and the major research fields represented within their corpora.
5 Discussion
Herein we outline our hierarchical density based tree construction method, applied onto LLM embeddings from large text corpora. Applied on the abstracts of scientific publications published by the American University of Beirut and TU Wien, our method produces interpretable, multi-resolution trees that reveal how topics, semantic fields, and research areas are organized relative to each other. By combining dimensionality reduction, adaptive multi-scale clustering, and large-context LLM-based annotation, we establish a pipeline capable of uncovering latent structure even in datasets lacking standardized or reliable taxonomies.
5.1 Insights from Benchmark Datasets
The benchmark datasets provide a reference for the applicability of our method.
Application of our method on the 20 newsgroups dataset revealed that the inferred tree was capable to recover intuitive relationships such as the proximity of sports-related topics, technical computing groups and overlaps between religious and political discussions. The hierarchical approach further uncovered that sports related discussions are further elongated from the different groups only merging to the whole tree close its root.
The AG news dataset similarly shows that World and Sports articles form well separated subtrees, but that Business and Science/Technology headlines exhibit a substantial overlap. This result can intuitively explained by the topical similarities between these two categories. Our tree based method as such creates an intrinsic semantic organization derived from LLM-embeddings rather then enforcing a predefined label boundary.
The French doctoral theses dataset provided a further test for our algorithm. This dataset included “discipline” annotations. However these are rather arbitrarily chosen and unstandarized. We transcribed these labels first into a binary annotation of either being part of natural sciences or humanities. To our surprise the theses abstracts of labeled humanities formed a clear subtree in our tree. Encouraged by these findings we progressed to transcribe the “discipline” labels into the OECD FOS annotation. We colored the tree according to it and found thanks to the hierarchical structure that engineering and medical sciences merge earlier than either social sciences or humanities. A relationship that as such cannot be derived from the FOS classification. At the same time we see that finer grained subtrees do correspond cleanly to the FOS subclasses, indicating that our inferred structure captures conceptual relationships that differ from established disciplinary boundaries.
These observations support the idea that our tree-based representations can serve as a complementary, data-driven alternative to traditional classification schemes.
The IMDB 50K Reviews dataset yields a very homogeneous tree, reflecting that the embedding models used in this study are not able to distinguish the semantic differences between the movie reviews in a meaningful way. The labels of this dataset are largely dispersed across the inferred tree, outlining the limitations of our approach, where the embeddings as outlined here do not encode as here sentiment or an other required topic separator.
5.2 Institution-Specific Research Landscapes
These datasets are at the core of this research as they illustrate that using our in here presented pipeline one can derive research classifications ab-initio, under the given that a recent, well trained LLM-embedding model exists. In both datasets, from the American University in Beirut and from TU Wien the inferred trees recover dominant research areas, internal substructures and relationships between disciplines without relying on a predefined taxonomy.
At AUB, the strong presence of medical and health-related research manifests as a large, cohesive subtree branching into public health, nutrition, epidemiology, and related fields. The emergence of a well-defined humanities subtree, in particular in Middle Eastern Studies and Conflict Resolution, highlights the institution’s regional and cultural focus. The structural proximity between engineering and humanities observed in the tree suggests interdisciplinary connections that may not be apparent from administrative classifications alone.
At TU Wien, the tree reflects the institution’s technical orientation, with prominent subtrees in materials science, quantum physics, information technology, and engineering. Despite the overall homogeneity of the corpus, the tree reveals meaningful internal organization. These findings demonstrate that the method can capture institution-specific research profiles and internal structure even in relatively focused corpora.
5.3 Methodological Considerations
Several methodological insights emerge from this study. First, dimensionality reduction via PCA plays a crucial role in enhancing tree structure. Using $L_{2}$ distances on a small number of principal components consistently yielded more articulated and interpretable trees than cosine distances in the full embedding space. This suggests that PCA acts as an effective feature-selection mechanism, suppressing noise and emphasizing the most informative semantic dimensions.
Second, the adaptive multi-$\epsilon$ DBSCAN procedure enables the recovery of hierarchical structure without imposing a fixed density threshold. By progressively relaxing density constraints, the method naturally captures cluster mergers across scales, producing a tree structure analogous to a dendrogram but grounded in density-based clustering.
Finally, while LLM-based annotation substantially improves interpretability, it also introduces limitations. Large clusters often receive overly general or unstable labels, reflecting current constraints of long-context inference. However, even imperfect annotations provide valuable guidance for navigating and interpreting complex tree structures.
5.4 Limitations and Future Directions
Inferred trees are inherently constrained by the representational capacity of the embedding models. Semantic dimensions that are weakly encoded, such as fine-grained sentiment or subtle disciplinary distinctions, cannot be reliably recovered through clustering alone. Future work could explore domain-specific or task-adapted embedding models to address this limitation.
In addition, annotation remains computationally expensive and sensitive to model variability. More robust strategies, such as hierarchical annotation, or summerization strategies tied together with annotation strategies may improve stability and scalability.
Looking beyond text-based corpora, a particularly promising direction is the extension of this framework to other data modalities. The method itself is agnostic to the origin of the embeddings and could, in principle, be applied to image embeddings, audio embeddings, or multimodal representations. For example, trees inferred from image embeddings could reveal visual themes or stylistic groupings, while trees built from music embeddings might uncover genre structure, rhythmic patterns, or harmonic similarities. Applying the proposed approach to such modalities would allow for the exploration of latent structure in domains where hierarchical organization is often implicit rather than explicitly labeled.
6 Conclusion
Overall, our results indicate that hierarchical density clustering of LLM embeddings provides a powerful and flexible framework for mapping semantic structure in large text corpora. The method reveals meaningful relationships that both complement and challenge existing classification systems. By enabling data-driven, multi-resolution representations of textual domains, this approach offers new opportunities for research assessment, knowledge mapping, and other exploratory analysis.
Acknowledgments
TH thanks the TU Wien Campus IT - Datalab - AI Team and EuroCC Austria for supporting this research. JB thanks the Center for Advanced Mathematical Sciences (CAMS) and the Artificial Intelligence, Data Science and Computing Hub (AI-DSC), for supporting this research.
We also thank the TU Wien Service Unit for Research Information Systems for providing the TU Wien dataset, the library of the American University of Beirut for providing the AUB dataset and Austrian Scientific Computing (ASC) for granting early access to the MUSICA supercomputing infrastructure:
https://asc.ac.at
Code and Data Availability
Code, TU Wien and AUB Datasets are available on github:
https://github.com/haschka/semantic-trees

### v2

@@
-Semantic Tree Inference on Text Corpa using a Nested Density Approach together with Large Language Model Embeddings
+Discovering Multi-Scale Semantic Structure in Text Corpora Using Density-Based Trees and LLM Embeddings
 Abstract
-Semantic text classification has undergone significant advances in recent years due to the rise of large language models (LLMs) and their high dimensional embeddings.
-While LLM-embeddings are frequently used to store and retrieve text by semantic similarity in vector databases, the global structure semantic relationships in text corpora often remains opaque.
-Herein we propose a nested density clustering approach, to infer hierarchical trees of semantically related texts.
-The method starts by identifying texts of strong semantic similarity as it searches for dense clusters in LLM embedding space.
-As the density criterion is gradually relaxed, these dense clusters merge into more diffuse clusters, until the whole dataset is represented by a single cluster - the root of the tree.
-By embedding dense clusters into increasingly diffuse ones, we construct a tree structure that captures hierarchical semantic relationships among texts.
-We outline how this approach can be used to classify textual data for abstracts of scientific abstracts as a case study.
-This enables the data-driven discovery research areas and their subfields without predefined categories.
-To evaluate the general applicability of the method, we further apply it to established benchmark datasets such as the 20 Newsgroups and IMDB 50k Movie Reviews, demonstrating its robustness across domains.
-Finally we discuss possible applications on scientometrics, topic evolution, highlighting how nested density trees can reveal semantic structure and evolution in textual datasets.
-1 Background
-Text classifications date back at least to bag of words models [bagofwords].
-In these early approaches, each text is represented as a vector of word frequencies, disregarding the order of words.
-However, word order is a key semantic feature as changing it can completely alter the meaning of a phrase.
-Subsequent refinements introduced term frequency and inverse document frequency (tf-idf) normalization [tfidf], which weights words by their relative importance within a corpus.
-Later, Latent Semantic Analysis (LSA) [lsa-first, lsa, lsa-intro] applied singlar value decomposition to capture latent relationships among words and documents, enabling more robust document clustering and topic modeling.
-Easier to scale and based on probabilistic properties instead of linear algebra Latent Dirichlet Allocation (LDA) [lda] and Hirachical Dirichlet Allocation (HDA) [hda] have overtaken the field of topic modelling.
-Neural network based approaches such as word2vec [word2vec] have arrived in the middle of the last decade.
-These models could learn the relationships between words in a semantic proximity sense, but still not in a grammatical sequence.
-The rise of the transformer model [transformer] and the popularization of LLMs allowed for the use of LLM embeddings for document classification, and document retrieval.
-The transformer model is superieur as it efficiently encodes and uses the position of the words in the text as it generates embeddings.
-LLM embeddings have a such a vast field of applications and [llm-embed-chatbot, llm-embed-clust] commonly found in retrieval augmented generation (RAG) pipelines [rag, llm-embed-rag].
-All of the aforementioned methods project text into a mathematical embedding space, where pairwise distances and clustering can be applied to reveal latent semantic structures.
-In parallel, hierarchical density clustering techniques have proven highly effective in fields such as biological species classification, phylogenetics [mnhn-tree-tools], and molecular dynamics, where they are used to analyze protein conformations and structural evolution [md-tree-tools].
-Here, we extend these hierarchical density clustering methods to textual embeddings, constructing tree-based representations of semantically related texts.
-This approach goes beyond flat clustering by revealing how topics, subjects, or emotions (depending on the embedding model) relate to one another within a coherent hierarchical framework.
+Recent advances in large language models enable documents to be represented as dense semantic embeddings, supporting similarity-based operations over large text collections.
+However, many web-scale systems still rely on flat clustering or predefined taxonomies, limiting insight into hierarchical topic relationships.
+In this paper we operationalize hierarchical density modeling on large language model embeddings in a way not previously explored.
+Instead of enforcing a fixed taxonomy or single clustering resolution, the method progressively relaxes local density constraints, revealing how compact semantic groups merge into broader thematic regions.
+The resulting tree encodes multi-scale semantic organization directly from data, making structural relationships between topics explicit.
+We evaluate the hierarchies on standard text benchmarks, showing that semantic alignment peaks at intermediate density levels and that abrupt transitions correspond to meaningful changes in semantic resolution.
+Beyond benchmarks, the approach is applied to large institutional and scientific corpora, exposing dominant fields, cross-disciplinary proximities, and emerging thematic clusters.
+By framing hierarchical structure as an emergent property of density in embedding spaces, this method provides an interpretable, multi-scale representation of semantic structure suitable for large, evolving text collections.
 1 Introduction
-Topic modeling and semantic clustering have long aimed to uncover the latent structure of large text corpora.
-With the rise of large language models (LLMs) and their embeddings, it has become possible to represent textual semantics with unprecedented precision.
-However, most embedding-based clustering approaches produce flat partitions of the semantic space, providing limited insight into how topics or concepts are hierarchically related or evolve across scales.
-In this work, we introduce a hierarchical density-based clustering approach that infers tree structures from LLM embedding spaces.
-The method builds upon previous applications of hierarchical density clustering in bioinformatics and molecular dynamics, where it has been successfully used to construct phylogenetic or conformational trees from complex, high-dimensional data [mnhn-tree-tools, md-tree-tools].
-By adapting this framework to text embeddings, we enable the reconstruction of semantic hierarchies.
-We show how dense clusters of highly similar documents merge into broader, more diffuse clusters as the similarity threshold is relaxed.
-To construct such hierarchical structures, we employ a nested density-based clustering procedure inspired by the DBSCAN algorithm [dbscan].
-The core idea is to identify compact, high-density regions of semantically similar embeddings and then progressively relax the density criterion to reveal broader, more diffuse groupings.
-This iterative relaxation enables clusters discovered at a high-density level to merge into larger clusters at lower density levels, naturally producing a hierarchical tree.
-Formally, clusters of semantically similar texts that that are densly connected in their respective embedding space are determined by the condition:
-$\rho_{L}>\frac{\mathrm{minpts}}{V(\epsilon_{L})},$ (1)
-where $\rho_{L}$ represents the denstiy at a given iteration $L$.
-The right hand side is given by starting parameters to the well known Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm [dbscan], with $\epsilon_{L}$ defining a radius at iteration $L$ and $\mathrm{minpts}$ a minimum number of points within this radius, effectively defining the density of a continious dense ensemble to be found by the algorithm.
-To explore the semantic structure at multiple scales, the radius $\epsilon_{L}$ is incrementally expanded by a small step $\Delta\epsilon$ in each iteration.
-This results in a sequence of density thresholds: As $\epsilon$ increases, the volume in equation (1) expands leading to lower densities and the identification of more diffuse clusters in each iteration.
-These new clusters encapsulate denser ones from previous iterations, allowing us to embed fine-grained semantic clusters into coarser ones, thereby forming a hierarchical tree of semantic relationships.
-For a subsequent layer (iteration) of the tree $L+1$ it follows, equation (2):
-$\rho_{L+1}>\frac{\mathrm{minpts}}{V(\epsilon_{L}+\Delta\epsilon)},$ (2)
-or alternatively $\epsilon_{L}=\epsilon_{L-1}+\Delta\epsilon$.
-In our previous work [mnhn-tree-tools], we demonstrated that using this algorithm allows us to construct a hierarchical tree structure by embedding dense clusters into progressively more diffuse clusters until all elements of the conformational space merge into a single cluster - the root of the tree.
-Figure 1 outlines the tree building procedure conceptually.
-On the left, a text embedding space is shown, where each colored sphere represents a piece of text.
-On the right the tree built by the algorithm, from the text embedding space on the left, is shown.
-One can observe how the grape, and orange colored texts fusion as the density criteria is relaxed into the blue ensemble, encompassing both clusters.
-At an intermediate step the green cluster stays separate, while the violet texts are not encompassed and seen as outliers.
-Finally, as the density criterion is further relaxed all texts are only part of one cluster, represented in gray on the right tree-outline.
-Figure 1: Tree building from a dataset represented by its embedding space.
-On the left, an embedding space with clusters of different densities is shown.
-One can see that the orange, and grape colored clusters embed into a larger blue ensemble.
-A conceptual tree built from such a two dimensional embedding space is shown on the right.
-Our algorithm allows an organization for instance to show how documents within the organization are related.
-It initially finds separated clusters of highly semantically similar documents (the outer leafs of the tree), that merge, layer by layer together to more diffuse clusters, until all documents form a single cluster (the root of the tree).
-We demonstrate our approach on several corpora, including collections of scientific abstracts and benchmark datasets such as 20 Newsgroups and IMDB 50K reviews.
-In these experiments, we show how the inferred semantic trees capture both fine-grained clusters of related documents and higher-level thematic organization.
-Applied to the publication records of institutions such as TU Wien and the American University of Beirut, our method reveals naturally emerging research areas and structural relationships between topics.
-Once a semantic tree is constructed, an important challenge lies in interpreting and annotating its clusters.
-We address this by leveraging LLMs themselves: by prompting a model with the text content of each cluster, we can automatically generate concise, human-readable summaries or titles for each node in the hierarchy.
-This creates an interpretable and fully automated pipeline for semantic structure discovery in large text corpora.
-3 Materials and Methods
+Understanding the semantic organization of large text corpora is a central challenge in knowledge representation, information retrieval, and semantic web research.
+As digital collections continue to grow in size and diversity, there is an increasing need for methods that not only group documents by similarity but also expose how concepts relate to one another across multiple levels of abstraction.
+Such structure is essential for tasks ranging from exploratory search and corpus navigation to the analysis of disciplinary landscapes and knowledge evolution.
+Recent advances in large language models (LLMs) have substantially improved the representation of textual meaning.
+High-dimensional embedding models encode semantic similarity with remarkable fidelity and have become a standard component of modern information systems, particularly in vector databases and semantic retrieval pipelines.
+While these embeddings enable effective similarity search and flat clustering, they typically provide limited insight into the global organization of semantic space.
+In particular, they do not directly reveal how fine-grained concepts combine into broader themes, nor how semantic resolution changes as one moves from local to global structure.
+Hierarchical representations offer a natural solution to this limitation.
+By organizing documents into nested structures, hierarchies can capture relationships between topics at multiple semantic scales.
+Traditional hierarchical approaches in text analysis, however, often rely on strong modeling assumptions, predefined taxonomies, or global distance matrices that are sensitive to noise and high dimensionality.
+As a result, they can struggle to scale to large corpora or to adapt flexibly to heterogeneous textual content.
+In this work, we propose a data-driven approach for reconstructing semantic hierarchies directly from text embedding spaces.
+Rather than imposing a predefined ontology or topic structure, we infer in an unsupervised manner hierarchical relationships by progressively revealing structure in the embedding space itself.
+The core idea is to identify groups of highly similar, documents and to examine how these groups merge as semantic similarity criteria are gradually relaxed.
+This process yields a tree-like structure in which terminal branches correspond to tightly related (dense) texts, while higher-level nodes represent increasingly (diffuse) general semantic groupings.
+A key advantage of this approach is that it produces a single, unified representation that simultaneously captures local semantic coherence and global thematic organization.
+The resulting trees are not restricted to a fixed number of clusters or levels and naturally accommodate documents that do not belong to any dense semantic group at finer resolutions.
+Such behavior is particularly important for real-world corpora, where topical boundaries are often fuzzy and unevenly populated.
+We demonstrate the utility of this framework on several benchmark datasets commonly used in text classification and topic analysis, including 20 Newsgroups, IMDB 50K reviews, and AG News.
+These experiments allow us to assess how well the inferred hierarchies align with established labels and to analyze how semantic agreement varies across hierarchical depth.
+Beyond benchmarks, we apply the method to large collections of scientific abstracts from academic institutions, where no single ground-truth taxonomy is assumed.
+In these cases, the inferred trees reveal institution-specific research profiles, cross-disciplinary connections, and emerging thematic areas.
+To support interpretability, we further integrate large language models as semantic annotators.
+By summarizing the textual content associated with individual tree nodes, we generate concise, human-readable descriptions of semantic regions within the hierarchy.
+This results in an end-to-end pipeline that combines embedding-based structure discovery with automated semantic labeling.
+Overall, this work contributes a flexible and interpretable method for uncovering hierarchical semantic structure in large text corpora.
+By emphasizing multi-scale organization, data-driven discovery, and semantic transparency, the proposed approach aligns naturally with the goals of semantic web research and provides new tools for exploring complex textual knowledge spaces.
+2 Related Work
+2.1 Text Representation for Semantic Analysis
+Text classification and organization have long relied on vector-based representations to capture semantic relationships within document collections.
+Early approaches such as bag-of-words models represent texts as unordered collections of terms, enabling efficient comparison but ignoring syntactic structure and semantic context [11].
+Subsequent developments introduced weighting schemes such as term frequency–inverse document frequency (tf–idf), which improved discrimination by accounting for term importance within a corpus [30].
+Latent Semantic Analysis (LSA) extended these ideas by projecting document-term matrices into lower-dimensional latent spaces using singular value decomposition, enabling the discovery of latent semantic associations [7, 10, 17].
+Probabilistic topic models, most notably Latent Dirichlet Allocation (LDA) and its hierarchical extensions, introduced generative frameworks for uncovering thematic structure in document collections [4, 31].
+While effective for identifying coarse-grained topics, these models require predefined assumptions about topic structure and often struggle with interpretability at fine semantic resolution [5].
+Neural embedding models, including word2vec and related architectures, marked a transition toward distributed semantic representations learned from large corpora [23].
+However, these approaches primarily capture word-level proximity and do not fully encode compositional or contextual meaning.
+The introduction of transformer-based architectures [32] and large language models (LLMs) enabled the generation of contextualized document-level embeddings that capture syntactic structure, semantics, and discourse-level information.
+LLM embeddings are now widely used in information retrieval, semantic search, and retrieval-augmented generation pipelines [18, 6].
+Despite their expressive power [9, 3, 2] their use in information organization is often limited to nearest-neighbor retrieval or flat clustering, leaving the global semantic structure of document collections largely unexplored.
+2.2 Clustering and Hierarchical Organization of Text
+Beyond representation, organizing documents into meaningful structures is a core challenge in information processing.
+Flat clustering methods such as k-means [20, 14] or DBSCAN [8] are frequently applied to document embeddings to identify groups of related texts.
+While computationally efficient, flat partitions impose a single resolution of analysis and obscure relationships between topics at different levels of semantic abstraction.
+Hierarchical approaches aim to overcome this limitation by explicitly modeling multi-level structure.
+Agglomerative hierarchical clustering constructs dendrograms from pairwise distances but requires global distance matrices and linkage criteria that are sensitive to noise, scaling, and high dimensionality.
+Similarly, hierarchical topic models rely on strong probabilistic assumptions and often exhibit limited robustness when applied to large, heterogeneous corpora [15, 25].
+Density-based clustering offers an alternative by identifying regions of high local similarity without assuming convex cluster shapes.
+While such methods have been successfully applied in spatial data analysis and other high-dimensional domains [13, 12], their use in text and information retrieval has largely focused on extracting flat clusterings rather than interpreting the full hierarchical structure.
+As a result, the potential of density-based hierarchies for revealing multi-scale semantic organization in text corpora remains to be explored.
+2.3 From Density-Based Clustering to Semantic Trees
+Hierarchical density-based clustering has been successfully applied in other high-dimensional domains, such as bioinformatics and molecular dynamics, to reconstruct phylogenetic relationships and conformational landscapes [13, 12].
+In these settings, trees are constructed by progressively relaxing density constraints, embedding compact structures into broader ones and yielding interpretable representations of multi-scale structure.
+This work adapts this paradigm to the organization of document embeddings.
+Rather than treating the hierarchy as an intermediate artifact or selecting a single optimal partition, we construct and analyze the full density-based tree as the primary representation of semantic structure.
+The resulting semantic tree captures how documents group and merge across similarity scales, enabling analysis at multiple levels of abstraction, from fine-grained topical distinctions to broad thematic domains.
+Crucially, the construction is independent of predefined taxonomies or label sets.
+This makes the approach well suited to exploratory analysis of large and heterogeneous corpora, where existing classifications are often incomplete, inconsistent, or entirely absent.
+2.4 Interpretability and Automatic Annotation
+Hierarchical organization alone is insufficient without interpretability.
+Large trees derived from embedding spaces require semantic summaries that allow users to understand and navigate the structure.
+Recent advances in LLMs enable large-context inference, making it possible to generate concise semantic descriptions for collections of documents.
+We leverage this capability to automatically annotate tree nodes by prompting LLMs with the textual content of each cluster.
+This produces human-readable labels that summarize the dominant themes represented at different levels of the hierarchy.
+The result is an end-to-end pipeline that combines unsupervised structure discovery with automatic semantic interpretation, addressing a long-standing challenge in large-scale document organization.
+3 Methodology
 3.1 Datasets
@@
   • 20 Newsgroups: The 20 Newsgroups dataset is a widely used benchmark corpus for text classification and topic modeling.
-It consists of approximately 20,000 newsgroup posts collected from 20 distinct discussion groups on topics such as science, politics, sports, religion, and technology [20newsgroups].
-  • IMDB 50K Reviews: A corpus of 50,000 movie reviews from the Internet Movie Database (IMDB), each labeled with a binary sentiment (positive or negative) [50kmovies].
-  • AG News: A large-scale news dataset consisting of short article titles and descriptions categorized into four topics: World, Sports, Business, and Science/Technology. [agnews].
+It consists of approximately 20,000 newsgroup posts collected from 20 distinct discussion groups on topics such as science, politics, sports, religion, and technology [24].
+  • IMDB 50K Reviews: A corpus of 50,000 movie reviews from the Internet Movie Database (IMDB), each labeled with a binary sentiment (positive or negative) [19].
+  • AG News: A large-scale news dataset consisting of short article titles and descriptions categorized into four topics: World, Sports, Business, and Science/Technology. [33].
   • Theses-fr-en: English abstracts of theses defended in France: The French government publishes by means of its open data initiative a dataset containing all theses defended in France since 1985.
@@
 Statistics about the different datasets are shown in table 1.
+Table 1: Statistics about the text corpa used.
+n(Texts) is the number of unique texts in the dataset, mean and stddev represent the mean and standard deviation of the number of words contained in each text of the dataset.
 n(Texts) mean stddev
+AUB 15982 237.73 98.24
 TU Wien 27506 168.85 73.79
-AUB 15982 237.73 98.24
 20newsgroups 11314 287.47 541.46
@@
 theses-fr-en 180939 261.89 112.70
-Table 1: Statistics about the text corpa used.
-n(Texts) is the number of unique texts in the dataset, mean and stddev represents the mean and standard deviation of the number of words contained in each text of the dataset.
 3.2 Embeddings
-Embeddings for the datasets outlined in section 3.1 were generated using either the Qwen3-Embedding-8B [qwen3embedding] or SFR-Embedding-Mistral [SFRembedding] model.
-These models were chosen selected based on their high performance in the MTEB-Leaderboard found on Huggingface, and the their practical deployability on the computing infrastructure accessible to us.
-4096 dimensonal embedding vectors obtained for each text from a dataset were stored in an in house developed vector database.
+Embeddings for the datasets outlined in section 3.1 were generated using either the Qwen3-Embedding-8B [34] or SFR-Embedding-Mistral [22] model.
+These models were chosen selected based on their high performance in the MTEB-Leaderboard found on Huggingface, and their practical deployability on the computing infrastructure accessible to us.
+4096-dimensonal embedding vectors obtained for each text from a dataset were stored in an in house developed vector database.
 Each text sample was transformed into a 4096-dimensional embedding vector.
@@
 3.3.2 Adaptive Density Clustering with DBSCAN
-For each dataset, we constructed trees using an adaptive density-based clustering procedure built upon successive runs of the DBSCAN algorithm.
-Each DBSCAN iteration identifies clusters at a specific density level, controlled by a triplet of parameters: the initial neighborhood radius $\epsilon_{0}$, the incremental radius step $\Delta\epsilon$, and the minimum number of points within the radius ($\mathrm{minpts}$).
-In each successive run, the radius is increased by $\Delta\epsilon$ while keeping $\mathrm{minpts}$ constant.
-This progressively lowers the effective density threshold and allows previously dense clusters to merge into more diffuse ones.
-Whenever the number of clusters found in a given iteration is smaller than in the preceding step, the newly identified clusters are stored, and the clusters from the prior iteration are embedded within them.
-This iterative embedding process builds the hierarchical tree structure described in Figure 1, where dense clusters at outer layers merge stepwise into broader, semantically diffuse clusters toward the root.
+Semantic clustering of text documents was performed using a density-based method adapted from DBSCAN [8].
+At iteration $L$, clusters were defined as regions where the local density $\rho_{L}$ satisfies
+$\rho_{L}>\frac{\mathrm{minpts}}{V(\epsilon_{L})},$ (1)
+where $\epsilon_{L}$ is the neighborhood radius defining a volume $V(\epsilon_{L})$, and $\mathrm{minpts}$ is the minimum number of points required within that volume.
+Unlike HDBSCAN [21], which constructs a hierarchy internally but typically outputs a single flat clustering, this method explicitly preserves hierarchical structure across successive density thresholds.
+To capture semantic structure at multiple scales, $\epsilon_{L}$ was increased in fixed increments $\Delta\epsilon$, generating progressively lower density thresholds.
+Clusters identified at each step contained denser clusters from previous iterations.
+For iteration $L+1$, the density criterion becomes
+$\rho_{L+1}>\frac{\mathrm{minpts}}{V(\epsilon_{L}+\Delta\epsilon)}.$ (2)
+For each dataset, hierarchical trees were constructed using successive runs of DBSCAN.
+Each iteration is controlled by three parameters: the initial radius $\epsilon_{0}$, the incremental step $\Delta\epsilon$, and the minimum points $\mathrm{minpts}$.
+In each successive run, the radius is increased while keeping $\mathrm{minpts}$ constant, progressively lowering the effective density threshold.
+Previously dense clusters merge into larger, more diffuse clusters, and whenever fewer clusters are found than in the preceding iteration, the new clusters are stored, embedding the previous ones within them.
+Figure 1 illustrates this process conceptually.
+Figure 1: Tree construction from a dataset represented in embedding space.
+Left: embedding space with clusters of different densities; each colored sphere represents a text document.
+Right: conceptual tree built from the embedding space, showing how denser clusters (orange and grape) merge into larger ensembles (blue) as the density threshold is relaxed, while other clusters (green) remain separate and some points (violet) are considered outliers.
+Eventually, all points merge into a single root cluster (gray).
+Unlike methods that optimize a predefined objective (e.g., topic coherence, likelihood, or cluster compactness), this approach constructs a hierarchical semantic representation directly from local density relations in the embedding space.
+It allows semantic groupings to emerge across multiple resolutions without assuming a fixed taxonomy, linkage criterion, or global distance structure.
+To our knowledge, this is the first method that operationalizes nested density relaxation as a general-purpose mechanism for semantic tree induction in text corpora.
 3.4 Tree Annotation
@@
 To generate these annotations, we employed a Large Language Model (LLM) capable of handling large textual contexts.
-Specifically, we used the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4] in 6-bit quantized form, which provides a nominal context window of up to 11 million tokens.
+Specifically, we used the Llama-4-Scout-17B-16E-Instruct model [1] in 6-bit quantized form, which provides a nominal context window of up to 11 million tokens.
 Due to technical constraints, our inference pipeline successfully operated with a maximum context size of approximately 600,000 tokens.
@@
 […]
-In the case were the cluster had to be sampled because it contained more than 1500 abstracts all resulting labels were subsequently fed into a second LLM pass.
-Again the gpt-oss-20b model [gpt-oss] in 8-bit quantized form is used in such a case where the model inferred the overarching research field encompassing all $s$ subfields.
+In the case where the cluster had to be sampled because it contained more than 1500 abstracts all resulting labels were subsequently fed into a second LLM pass.
+Again the gpt-oss-20b model [28] in 8-bit quantized form is used in such a case where the model inferred the overarching research field encompassing all $s$ subfields.
 An overview about the different paths that annotation might undergo is outlined in Figure 2.
-Figure 2: Shematic represantation of the tree node / cluster annotation process
+Figure 2: Schematic representation of the tree node / cluster annotation process.
 Tree annotation is computationally intensive.
 The AUB tree required annotating 33,862 text files, while the TU Wien tree required 62,012, including oversampling.
-Inference was performed using the llama.cpp [llama-cpp] engine on the MUSICA supercomputer (Austria).
+Inference was performed using the llama.cpp [26] engine on the MUSICA supercomputer (Austria).
 Up to 80 NVIDIA H100 GPUs were employed concurrently during annotation.
@@
 Multi-node inference, which could allow for a higher context token limit, was not attempted as MUSICA is currently in its test phase; instead, parallelism was achieved by distributing text files across multiple nodes for simultaneous inference.
+3.5 Evaluation Metrics for Hierarchical, non-Consistent Partitions
+To quantitatively assess how well the inferred semantic trees align with existing labels in benchmark datasets, we employ a combination of external clustering validation measures: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI).
+For the evalution of these metrics we make use of their implementation in Scikit-Learn [29].
+These metrics are applied independently at each tree layer, treating the clustering at that layer as a flat partition of the dataset.
+ARI and NMI are widely used measures for comparing a predicted clustering against a ground-truth labeling.
+ARI evaluates agreement between two partitions based on pairwise co-assignment, adjusted for chance, while NMI measures the shared information between cluster assignments and class labels, normalized to account for differing numbers of clusters.
+In our setting, these metrics are computed per tree layer using only the documents assigned to clusters at that layer.
+This is necessary because the density-based procedure may leave a fraction of documents unclustered at higher density levels, leading to non-exhaustive and non-consistent partitions across layers.
+While ARI and NMI are not inherently designed for hierarchical or incomplete clusterings, they remain informative for assessing local alignment between inferred clusters and known labels where such overlap exists.
 3.6 Label Transcription
@@
 In a first test we cross annotated these unique discipline labels into a binary classification of either humanities or natural sciences.
-We then proceeded to obtain a more coherent and standardized annotation, we used the gpt-oss-20b model to classify all 13,599 labels according to the OECD Fields of Science (FOS) 2002 classification [oecd-fos].
+We then proceeded to obtain a more coherent and standardized annotation, we used the gpt-oss-20b model to classify all 13,599 labels according to the OECD Fields of Science (FOS) 2002 classification [27].
 The resulting transcribed labels were then used to color the trees inferred from this dataset, allowing us to assess whether specific scientific fields tend to separate into distinct and meaningful branches.
@@
 Similarly, for the TU Wien dataset, coloration based on existing research classifications allowed us to assess how well our hierarchical density clustering recovers semantically coherent domains.
-Tree coloration was performed using utilities provided by MNHN-Tree-Tools [mnhn-tree-tools].
+Tree coloration was performed using utilities provided by MNHN-Tree-Tools [13].
 In particular, we employed the tree_map_for_split_set tool to map predefined clusters or labeled subsets onto the corresponding nodes of the hierarchical tree.
-The resulting color maps were generated for visualization using the newick-utilities package [newick-utilities], a software suite for the manipulation and visualization of phylogenetic trees.
+The resulting color maps were generated for visualization using the newick-utilities package [16], a software suite for the manipulation and visualization of phylogenetic trees.
 This approach allowed us to produce colored tree visualizations for each dataset.
@@
 We suggest that this effect is due to the inherent ability of PCA to yield principal components that maximize variance and that this feature selection process has a positive effect on our clustering algorithm.
-A similar effect was noted previously in the algorithmic supplement of [mnhn-tree-tools].
+A similar effect was noted previously in the algorithmic supplement of [13].
 A comparison between trees obtained with cosine distance and with the $L_{2}$-norm applied to PCA-reduced embeddings is shown in Figure 3.
@@
 The parameter settings were as follows: for cosine distance, initial $\epsilon=0.275$, $\Delta\epsilon=0.001$, and $\mathrm{minpts}=5$; for the $L_{2}$-norm on 10 principal components, initial $\epsilon=0.091$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$; and for the $L_{2}$-norm on 2 principal components, initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
-The trees shown in 3 were built from embeddings obtained form the Qwen3-Embedding-8B model.
-Figure 3: A comparison of trees built from the 20 Newsgroups dataset under different distance measures
+The trees shown in 3 were built from embeddings obtained from the Qwen3-Embedding-8B model.
+Figure 3: A comparison of trees built from the 20 Newsgroups dataset under different distance measures.
 4.2 Detection of Topics and Semantics
-Depending on the dataset, we applied different evaluation strategies to assess whether, and to what extent, the method proposed in this work produces meaningful and reliable results, given the computational constraints of our setup.
-The following subsections outline these evaluations and illustrate the behavior of our hierarchical density-based tree construction across diverse text corpora.
+To evaluate whether the proposed hierarchical density-based method recovers meaningful semantic structure, we combine qualitative analysis of inferred trees with quantitative evaluation against available ground truth labels.
+Since the output of our method is a hierarchy rather than a single flat partition, evaluation must account for the fact that clusters emerge, merge, and disappear across tree layers.
+For datasets with reference labels, we therefore project the hierarchy onto a sequence of flat clusterings, one per tree layer, and assess their agreement with the ground truth using the Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI).
+ARI measures the agreement between two partitions while correcting for chance, whereas NMI quantifies the amount of shared information between cluster assignments and labels, normalized to allow comparison across different numbers of clusters.
+Both measures are well established in information retrieval and clustering evaluation and are complementary: ARI is sensitive to exact cluster membership, while NMI captures coarser label alignment even when cluster boundaries are imperfect.
+Across the benchmark datasets, the evolution of ARI and NMI as a function of tree depth reveals dataset-specific behavior that reflects differences in semantic structure rather than methodological artifacts.
+For the 20 Newsgroups and AG News corpora, agreement with ground-truth labels increases from the finest partitions toward intermediate density levels, reaches a maximum, and then exhibits sharp decreases followed by extended plateaus as clusters merge at coarser scales.
+This pattern indicates that semantic alignment with predefined categories is strongest at intermediate resolutions, where fine-grained lexical variation has been absorbed while broader thematic distinctions are still preserved.
+Details are outlined in Figure 7.
+In contrast, the IMDB 50K Reviews dataset exhibits a different trend.
+Here, the highest ARI and NMI values occur closer to the leaves of the tree, within relatively fine-grained branches rather than near the root as shown in Figure 7.
+As density constraints are relaxed and clusters merge, agreement with sentiment labels decreases steadily, and no dominant peak is observed at intermediate or coarse levels.
+This behavior reflects the largely homogeneous semantic structure of the corpus, where sentiment is encoded as a subtle, distributed signal rather than as a dominant organizing principle.
+Consequently, sentiment-aligned clusters emerge only locally and do not persist as stable large-scale structures.
+Taken together, these results demonstrate that the proposed hierarchical density-based trees adapt to the intrinsic semantic organization of each dataset.
+Rather than enforcing a fixed optimal resolution, the method exposes multiple meaningful semantic scales, allowing the analyst to identify the levels at which different types of structure-topical, thematic, or affective, are most pronounced.
+Importantly, at high-density (outer) layers, only compact and highly coherent clusters are present, while diffuse documents remain unassigned.
+As density constraints are relaxed, clusters merge and eventually encompass the full dataset.
+ARI and NMI are therefore evaluated as a function of tree depth, allowing us to quantify how semantic agreement evolves across scales rather than at a single, arbitrarily chosen resolution.
 4.2.1 20 Newsgroups
-To evaluate whether our tree construction method recovers meaningful semantic structure, we applied it to the 20 Newsgroups dataset.
-The resulting trees were colored according to posting labels, as shown in Figure 4.
-For each newsgroup we generated a separate color map: posts from the target newsgroup are shown in blue, and posts from all other newsgroups in orange.
-Both color intensities follow a logarithmic scale proportional to the number of posts in each cluster, with white indicating the absence of posts of a given type.
-The inferred tree, as highlighted in figure 4, recovers many interesting relationships between the newsgroups:
+We first apply our method to the 20 Newsgroups dataset, which provides a well-studied benchmark with twenty topical labels.
+Figure 4 shows the inferred tree, colored according to individual newsgroup labels.
+Quantitatively, ARI and NMI peak at intermediate tree depths, indicating that the strongest alignment between inferred clusters and newsgroup labels occurs neither at the finest nor at the coarsest resolution.
+At very high densities, clusters are too fragmented to recover complete topics, while at low densities, semantically distinct groups merge into broader themes.
+This behavior is expected for a hierarchical method and confirms that meaningful topic structure emerges at specific semantic scales.
+Qualitatively, the tree recovers intuitive relationships between newsgroups:
   • Sports-related groups.
@@
 This intersection reflects the topical proximity and overlapping discussions between religion and politics within the dataset.
-  • Computer-related groups.
-The comp.* groups, including comp.sys.mac.hardware, comp.windows.x, comp.graphics, and comp.os.ms-windows, form a coherent subtree corresponding to computer hardware and software discussions.
-Some overlap with sci.electronics is visible, consistent with semantically similar content.
+  • Computer-related groups. The comp.* groups, including
+    comp.sys.mac.hardware, comp.windows.x,
+    comp.graphics, and
+    comp.os.ms-windows, form a coherent subtree corresponding to computer hardware and software discussions. Some overlap with sci.electronics is visible, consistent with semantically similar content.
   • Marketplace group.
@@
 This separation may stem from the unique vocabulary and specialized discourse typical of cryptography discussions.
-Overall the structure seems to intuitively make sense, outlining intervoven fields of discussion.
-Importantly, our toolchain does not merely cluster newsgroups into isolated groups; it also conveys the relative proximity and semantic distance between them.
-The resulting tree therefore serves as a compact visual instrument that captures both fine-grained cluster structure and the broader topology of topic similarity within the dataset, all within a single interpretable representation.
+These qualitative observations align with the quantitative ARI/NMI profiles: tree regions that correspond to stable semantic groupings also contribute most strongly to the metric peaks.
+Crucially, the tree representation conveys not only cluster membership but also relative semantic proximity between topics, which is not captured by flat clustering alone.
 Figure 4: Representations of the tree obtained from the 20 Newsgroups dataset.
-Trees were built from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] embeddings.
-Parameters used to build tree were: initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
+Trees were built from the first two principal components obtained from Qwen3-Embedding-8B [34] embeddings.
+Parameters used to build the tree were: initial $\epsilon=0.0039$, $\Delta\epsilon=0.00001$, $\mathrm{minpts}=5$.
 4.2.2 IMDB 50K Reviews
-The IMDB 50K Reviews dataset was used to evaluate whether our tree construction method, together with the underlying embedding models, is able to separate reviews by emotional polarity.
-We therefore built a tree from this dataset following the same procedure as for the other corpora.
-The resulting tree was then colored analogously to the 20 Newsgroups experiments: clusters containing predominantly positive reviews were assigned a blue shading, while clusters containing predominantly negative reviews were shaded in orange.
+We next evaluate the IMDB 50K Reviews dataset, which provides binary sentiment labels (positive vs.
+negative).
+Unlike topical datasets, this corpus is semantically homogeneous, and sentiment is expressed largely through subtle linguistic cues.
+ARI and NMI values remain comparatively low across all tree depths, with only a modest peaks at layers still in the high dense leaf node range.
+This indicates that the embedding models used encode sentiment only weakly relative to other semantic dimensions.
+The inferred tree, shown in Figure 5, reflects this quantitatively observed behavior: the structure is shallow, with few large subtrees and only localized regions dominated by either positive or negative reviews.
 As before, color intensity follows a logarithmic scale, with white indicating the absence of reviews of a given label and deeper blue or orange indicating higher concentration.
 Overlaps appear in darker, greenish to black tones.
-The resulting tree is shown in Figure 5.
-Compared to the other datasets, the IMDB tree appears noticeably less structured, suggesting that the underlying corpus is more homogeneous and lacks strong semantic subdivisions beyond sentiment.
-Nonetheless, we observe a few small subtrees near the periphery of the tree in which positive reviews form a clear majority.
-Additional localized regions with a majority of positive or negative reviews can also be identified.
-We speculate that these structures arise from repetitive linguistic patterns that occur more frequently in either positive or negative reviews.
-However, they do not form large coherent branches, and overall the dataset does not separate cleanly by sentiment.
-This observation suggests that the embedding models used: Qwen3-Embedding-8B [qwen3embedding] and SFR-Embedding-Mistral [SFRembedding] encode emotional polarity only to a limited extent.
+Small terminal branches enriched in positive or negative sentiment do exist, suggesting that repetitive phrasing or stylistic patterns occasionally induce local clustering.
+However, no large-scale sentiment-based separation emerges.
+These findings highlight an important methodological point: hierarchical semantic trees faithfully reflect the information encoded in the embedding space and should not be expected to recover distinctions that are not strongly represented by the underlying model.
+This observation suggests that the embedding models used: Qwen3-Embedding-8B [34] and SFR-Embedding-Mistral [22] encode emotional polarity only to a limited extent.
 The results highlight the importance of choosing embedding models appropriate for the expected semantic structure when constructing trees.
-We also attempted to annotate the IMDB tree with the Llama-4-Scout-17B-16E-Instruct model [meta2024llama4].
+We also attempted to annotate the IMDB tree with the Llama-4-Scout-17B-16E-Instruct model [1].
 The goal was to obtain genre-level descriptions of tree regions, under the assumption that certain movie genres might correlate with predominantly positive or negative reviews.
@@
 As such, deriving movie genres from this dataset seems infeasible with current methods.
-Figure 5: The IMDB 50K Reviews dataset colored by sentiments: The tree was built using the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings.
+Figure 5: The IMDB 50K Reviews dataset colored by sentiments: The tree was built using the first two principal components derived from SFR-Embedding-Mistral [22] embeddings.
 Parameters yield to the tree building algorithm were: initial $\epsilon=0.5$, $\Delta\epsilon=0.0001$, $\mathrm{minpts}=5$.
@@
 Figure 6: The tree obtained from the AG News dataset in different colorations.
-The tree shown was inferred from the first two principal components derived from SFR-Embedding-Mistral [SFRembedding] embeddings.
+The tree shown was inferred from the first two principal components derived from SFR-Embedding-Mistral [22] embeddings.
 Parameters yield to the tree building algorithm were: initial $\epsilon=0.225$, $\Delta\epsilon=0.0005$, and $\mathrm{minpts}=5$
 We observe that the world and sports categories form clearly separable structures: both branch off early from the root of the tree, resulting in two well-defined subtrees with only minimal contamination from other labels.
+This is also reflected by stronger ARI and NMI value peaks than in the other datasets as outlined in Figure 7.
 Around the root, however, a small number of scattered articles appear, suggesting that certain headlines are not easily classifiable by the embedding models.
@@
 We attribute this to the general type of embedding models that we are using here and that business headlines frequently mention science and technology and vice versa in this dataset.
+Figure 7: Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) as a function of tree depth for the 20 Newsgroups, IMDB 50K Reviews and AG News dataset.
+20 Newsgroups and AG News peak at intermittent clusters indicating that semantic agreement with ground-truth labels is maximized at intermediate density levels rather than at the finest or coarsest partitions.
 4.2.4 The French Theses Dataset: Theses-fr-en
@@
 Therefore, as outlined in Section 3.6, we first performed a cross-annotation step to obtain more tangible labels.
+As these labels were generated by us we did not apply ARI and NMI evaluations on this tree.
 As an initial experiment, we mapped the original annotations to a binary classification: humanities versus natural sciences.
 The resulting tree clearly separated humanities theses from the remainder of the corpus, forming a large, coherent branch as shown in Figure 8A.
-Encouraged by this result, we proceeded to apply a more detailed transcription based on the OECD Fields of Science and Technology (FOS) 2002 classification scheme [oecd-fos].
-In the tree three major branches are easily identifyable.
+Encouraged by this result, we proceeded to apply a more detailed transcription based on the OECD Fields of Science and Technology (FOS) 2002 classification scheme [27].
+In the tree three major branches are easily identifiable.
 The branch identified as “humanities” in the binary test contains predominantly Social Sciences (5) and Humanities (6) according to the FOS taxonomy.
@@
 This observation directly motivates the subsequent section of the paper.
-Figure 7: Tree generated from the French Theses theses-fr-en dataset using the first two principal components obtained from the embeddings yield by the SFR-Embedding-Mistral [SFRembedding] model:
+Figure 8: Tree generated from the French Theses theses-fr-en dataset using the first two principal components obtained from the embeddings yield by the SFR-Embedding-Mistral [22] model:
 A: Coloration according to binary classification Humanities and Natural Sciences.
@@
 In contrast, the institutional datasets are used to examine research structures ab initio: rather than testing whether a predefined taxonomy is recovered, we investigate how the tree itself organizes the institutional research landscape.
-Although the embedding models were taken trained on external corpora, and not further fine tuned, the resulting tree structure is independent of any curated ontology.
+Consequently, ARI and NMI are not applicable.
+Although the embedding models were trained, by their creators, on external corpora, and not further fine tuned, the resulting tree structure is independent of any curated ontology.
 This allows us to explore institution-specific disciplinary profiles, structural proximities between fields, and potential emerging clusters or cross-disciplinary connections.
 The results demonstrate that our method can capture meaningful organizational patterns without relying on predefined categories.
-  • At the American University of Beirut, medicine and health-related disciplines dominate the research corpus, reflecting the strong presence of the university hospital.
+  • American University of Beirut (AUB)
+    At AUB, medicine and health-related disciplines dominate the research corpus, reflecting the strong presence of the university hospital.
 This dominance is visible in the tree structure, where the entire upper half of the circular tree forms the largest and most cohesive subtree.
@@
     A distinct humanities subtree is also present, characterized by numerous subbranches labeled as Middle Eastern Studies and Conflict Resolution.
-In addition, an engineering subtree that covers engineering, physics, and mathematics, is clearly visible.
+In addition, an engineering subtree that covers engineering, physics, and mathematics is clearly visible.
 Notably, this engineering subtree merges with the humanities subtree at a higher density than with the medical subtree, suggesting a closer structural relationship between engineering and humanities research topics within this corpus than might be expected, which might be characteristic of the university’s liberal arts tradition.
@@
 As with the AUB tree, color intensity follows a logarithmic scale from white (no articles) to black (all articles in the dataset).
-Figure 8: Tree generated from the AUB dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings.
+Figure 9: Tree generated from the AUB dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [34] model embeddings.
 The tree was built with the following parameters: initial $\epsilon=0.0045$, $\Delta\epsilon=0.00001$ and $\mathrm{minpts}=5$.
-Figure 9: Tree generated from the TU Wien dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [qwen3embedding] model embeddings.
+Figure 10: Tree generated from the TU Wien dataset: The tree shown was inferred from the first two principal components obtained from Qwen3-Embedding-8B [34] model embeddings.
 The tree was built with the following parameters: initial $\epsilon=0.0028$, $\Delta\epsilon=0.000001$ and $\mathrm{minpts}=5$.
@@
 The AG news dataset similarly shows that World and Sports articles form well separated subtrees, but that Business and Science/Technology headlines exhibit a substantial overlap.
-This result can intuitively explained by the topical similarities between these two categories.
-Our tree based method as such creates an intrinsic semantic organization derived from LLM-embeddings rather then enforcing a predefined label boundary.
+This result can intuitively be explained by the topical similarities between these two categories.
+Our tree based method as such creates an intrinsic semantic organization derived from LLM-embeddings rather than enforcing a predefined label boundary.
 The French doctoral theses dataset provided a further test for our algorithm.
 This dataset included “discipline” annotations.
-However these are rather arbitrarily chosen and unstandarized.
+However these are rather arbitrarily chosen and unstandardized.
 We transcribed these labels first into a binary annotation of either being part of natural sciences or humanities.
@@
 The IMDB 50K Reviews dataset yields a very homogeneous tree, reflecting that the embedding models used in this study are not able to distinguish the semantic differences between the movie reviews in a meaningful way.
-The labels of this dataset are largely dispersed across the inferred tree, outlining the limitations of our approach, where the embeddings as outlined here do not encode as here sentiment or an other required topic separator.
+The labels of this dataset are largely dispersed across the inferred tree, outlining the limitations of our approach, where the embeddings as outlined here do not encode as here sentiment or any other required topic separator.
 5.2 Institution-Specific Research Landscapes
@@
 In addition, annotation remains computationally expensive and sensitive to model variability.
-More robust strategies, such as hierarchical annotation, or summerization strategies tied together with annotation strategies may improve stability and scalability.
+More robust strategies, such as hierarchical annotation, or summarization strategies tied together with annotation strategies may improve stability and scalability.
 Looking beyond text-based corpora, a particularly promising direction is the extension of this framework to other data modalities.